# Approximation and FPT Algorithms for Finding DM-Irreducible Spanning Subgraphs

### v2

@@ Abstract
Finding a minimum-weight strongly connected spanning subgraph of an edge-weighted directed graph is equivalent to the weighted version of the well-known strong connectivity augmentation problem, and it is NP-hard. For this problem, a simple $2$-approximation algorithm was proposed by Frederickson and Jájá (1981); surprisingly, it still achieves the best-known approximation ratio in general. Also, the unweighted problem was shown to be FPT (fixed-parameter tractable) by Bang-Jensen and Yeo (2008), if it is parameterized by the difference from the trivial upper bound of the optimal value. In this paper, we consider a generalized problem related to the Dulmage–Mendelsohn decompositions of bipartite graphs instead of the strong connectivity of directed graphs, and extend these approximation and FPT results to the generalized setting.
@@ 1 Introduction
A directed graph is called strongly connected if there exists a path between any pair of vertices. The strong connectivity augmentation problem (SCA) is, given a directed graph and the weights of additional edges, to find a minimum-weight subset of additional edges making the graph strongly connected. While SCA is NP-hard in general, it can be solved in linear time in the unweighted case (when any edge can be added and has the same weight) [9].
@@
The following problem, called the strongly connected spanning subgraph problem (SCSS), is polynomial-time equivalent to SCA in general: given a strongly connected directed graph with edge weight, find a minimum-weight spanning subgraph that is strongly connected. Frederickson and Jájá [14] gave a $2$-approximation algorithm for SCSS; this is based on a very simple observation (see Section 3.1) and, surprisingly, it still achieves the best known approximation ratio in general. For the unweighted case (when all the edges have the same weight, which is already NP-hard in contrast to SCA), several algorithms with better approximation ratios [18, 26, 27, 17] have been proposed. Also, Bang-Jensen and Yeo [2] showed that the unweighted case is FPT if it is parameterized by the difference from the trivial upper bound of the optimal value (see Sections 2.2 and 4.1); here, the problem is called FPT (fixed-parameter tractable) if it can be solved in $f(k)\cdot n^{O(1)}$ time, where $k$ and $f$ are the prescribed parameter and computable function, respectively, and $n$ is the input size.
@@
SCSS (or SCA) is generalized to several directions. One natural generalization is the $k$(-edge)-connected spanning subgraph (or $k$(-edge)-connectivity augmentation) problem, where a directed graph is called $k$-connected (or $k$-edge-connected) if there exists $k$ openly disjoint (or edge-disjoint) paths between any pair of vertices. The case when $k=1$ corresponds to SCSS (or SCA), and the general case is also well-studied [10, 12, 19, 24, 1].
@@
Bérczi, Iwata, Kato, and Yamaguchi [3] recently investigated another generalization of SCA, which is related to the Dulmage–Mendelsohn decomposition (the DM-decomposition) [7, 6] of a bipartite graph. The DM-decomposition gives a partition of the vertex set that reflects the structure of all maximum matchings therein, and a bipartite graph is DM-irreducible if its DM-decomposition consists of a single component. This property is closed under adding edges, and we analogously formulate the DM-irreducible spanning subgraph problem (DMISS): given a DM-irreducible bipartite graph with edge weight, find a minimum-weight spanning subgraph that is DM-irreducible. Then, DMISS indeed generalizes SCSS (cf. [3, Section 2.3]), and in terms of DMISS, Bérczi, et al. [3] proposed polynomial-time algorithms for the case when the input bipartite graph is unbalanced with the aid of weighted matroid intersection, and for the case when the input graph is a balanced complete bipartite graph and the weight of each edge is 00 or $1$, where a bipartite graph is balanced if the two color classes have the same number of vertices. In this paper, by extending the approaches to SCSS by Frederickson and Jájá [14] and by Bang-Jensen and Yeo [2], we propose a $2$-approximation algorithm for DMISS in general and an FPT algorithm for the unweighted case of DMISS.
@@
The rest of this paper is organized as follows. In Section 2, we describe necessary definitions, and formally state the problems and our results. In Section 3, we give a $2$-approximation algorithm for DMISS. In Section 4, we prove the unweighted case of DMISS is FPT parameterized by the difference from the trivial upper bound of the optimal value.
@@ 2.1 Definitions
For a graph $G$, we denote by $V(G)$ and $E(G)$ the vertex set and the edge set of $G$, respectively. A subgraph of $G$ is a graph $G^{\prime}$ such that $V(G^{\prime})\subseteq V(G)$, $E(G^{\prime})\subseteq E(G)$, and each edge $e\in E(G^{\prime})$ is between two vertices in $V(G^{\prime})$. A subgraph $G^{\prime}$ of $G$ is induced by a vertex set $X\subseteq V(G)$, denoted by $G[X]$, if $V(G^{\prime})=X$ and $E(G^{\prime})$ is the set of edges in $E(G)$ whose end vertices are both in $X$. We also define $G-X\coloneqq G[V\setminus X]$ for $X\subseteq V$. A subgraph $G^{\prime}$ of $G$ is spanning if $V(G^{\prime})=V(G)$. We often do not distinguish between a spanning subgraph $G^{\prime}$ and its edge set $E(G^{\prime})$.
A stable set in an undirected graph $G$ is a vertex set $X\subseteq V(G)$ such that no edge in $E(G)$ has its both end vertices in $X$. An undirected graph $G=(V,E)$ is bipartite if there exists a bipartition $(V^{+},V^{-})$ of $V$ such that $V^{+}\cap V^{-}=\emptyset$, $V^{+}\cup V^{-}=V$, and both $V^{+}$ and $V^{-}$ are stable sets in $G$ (possibly $V^{+}$ or $V^{-}$ is empty). To specify the bipartition, we denote by $G=(V^{+},V^{-};E)$. For a vertex set $X$ on one side (i.e., $X\subseteq V^{+}$ or $X\subseteq V^{-}$), let $\Gamma_{G}(X)$ denote the neighborhood of $X$, i.e., the set of vertices adjacent to some vertex in $X$. A bipartite graph $G=(V^{+},V^{-};E)$ is balanced if $|V^{+}|=|V^{-}|$.
@@
The DM-decomposition [7, 6] of a bipartite graph is formally defined by the distributive lattice formed by the minimizers of a submodular function, and a bipartite graph is called DM-irreducible if its DM-decomposition consists of a single component. We here employ its characterization as an alternative “definition” (cf. [3, Section 2.2]). [^1] The equivalent condition is also called elementary, e.g., in [23, 11].
Lemma 2.1 (cf. [11, Theorem 2.4.18]).
@@
    $G$ has a perfect matching, and for any perfect matching $M$ in $G$, $G_{M}$ is strongly connected.
Suppose that a graph $G$ is associated with edge weight $w\colon E\to\mathbb{R}$. We define the weight of each edge set $F\subseteq E$ as $w(F)\coloneq\sum_{e\in F}w(e)$.
@@ 2.2 Problems and Our Results
We formally state the problems dealt with in this paper. Recall that DMISS for the unbalanced bipartite graphs is tractable [3], and it is not considered in this paper.
@@
A strongly connected digraph $G=(V,E)$ with edge weight $w\colon E\to\mathbb{R}_{\geq 0}$.
@@
A DM-irreducible balanced bipartite graph $G=(V^{+},V^{-};E)$ with edge weight $w\colon E\to\mathbb{R}_{\geq 0}$.
@@
It is easy to observe that any inclusion-wise minimal strongly connected spanning subgraph of a digraph on $n$ vertices consists of at most $22n-$ 2 edges (by Lemma 3.1). Similarly, it is easy to observe that any inclusion-wise minimal DM-irreducible spanning subgraph of a balanced bipartite graph on $2n$ vertices ($n$ vertices on each side) consists of at most $23n-$ 2 edges (by Lemma 3.2). The following problems are the unweighted version of the above two problems parameterized by the difference from these trivial upper bounds (note that it is obvious from the definition that $|V|$ edges are necessary for either problem).
@@
A strongly connected digraph $G=(V,E)$.
@@
Test whether $G$ has a strongly connected spanning subgraph with at most $2|V|-2-k$ edges or not.
@@
A DM-irreducible balanced bipartite graph $G=(V^{+},V^{-};E)$.
@@
Test whether $G$ has a DM-irreducible spanning subgraph with at most $3|V^{+}|-2-k$ edges or not.
The results shown in this paper are summarized as follows, where $n=|V^{+}|=|V^{-}|$.
@@ 3 Approximation Algorithms for DMISS
In Section 3.1, as a base of our result, we describe a $2$-approximation algorithm for SCSS by Frederickson and Jájá [14]. In Section 3.2, we show a $3$-approximation algorithm for DMISS obtained by a naive extension of the result of Frederickson and Jájá. In Section 3.3, we improve it to a $2$-approximation algorithm with the aid of matroid intersection.
@@
The algorithm is based on the following characterization of strongly connected digraphs.
@@
For the input $(G=(V,E),w)$ of SCSS, the algorithm arbitrarily chooses a vertex $r\in V$ and finds a minimum-weight spanning in-arborescence $A_{\mathrm{in}}$ and out-arborescence $A_{\mathrm{out}}$ rooted at $r$; the existence of $A_{\mathrm{in}}$ and $A_{\mathrm{out}}$ follows from Lemma 3.1 $(1\Rightarrow 3)$. This can be done in $O(m+n\log n)$ time [15], where $n=|V|$ and $m=|E|$. By Lemma 3.1 $(2\Rightarrow 1)$, the union of $A_{\mathrm{in}}$ and $A_{\mathrm{out}}$ results in a strongly connected spanning subgraph of $G$, and the algorithm outputs this. Since any optimal solution has a spanning in-arborescence and out-arborescence rooted at $r$ by Lemma 3.1 $(1\Rightarrow 3)$ and their weights are at least those of $A_{\mathrm{in}}$ and $A_{\mathrm{out}}$, respectively, the weight of the output solution is at most twice the optimal value.
@@
For the first task, polynomial-time exact algorithms initiated by the Hungarian method [22] are well-known. For the second task, we apply the $2$-approximation algorithm described in the previous section. We then output the union of the solutions (for the second solution, we take the set of corresponding edges in $G$). This intuitively leads to a $3$-approximation algorithm for DMISS (as $1+2=3$), which can be formally shown as follows.
@@
We first give an easy but important observation. In short, for any $r\in V^{+}$, all the matching edges are included in every spanning $r$-in-arborescence in the auxiliary graph.
@@
    $T$ has a perfect matching $M$ in $G$ and $\overrightarrow{T\setminus M}\cup\overleftarrow{M}$ is a spanning in-arborescence of $G_{M}$.
@@
Let us call a spanning tree satisfying Statement 2 in Lemma 3.5 strongly balanced, and say that the vertex $r$ is the root and the tree is rooted at $r$. By Lemma 3.2, every DM-irreducible graph has a strongly balanced spanning tree, and then we formulate the first task as follows.
@@
A DM-irreducible bipartite graph $G=(V^{+},V^{-};E)$ with edge weight $w\colon E\to\mathbb{R}_{\geq 0}$.
@@
For the sake of simplicity of the following discussion, we consider the following problem, which is polynomially equivalent to the above problem and actually enough for our purpose.
Problem (Strongly Balanced Spanning Tree with Root (SBSTR)).
@@
A DM-irreducible bipartite graph $G=(V^{+},V^{-};E)$ with edge weight $w\colon E\to\mathbb{R}_{\geq 0}$, and $r\in V^{+}$.
@@
As we already mentioned, the second task (finding a minimum-weight spanning arborescence) can be done in $O(m+n\log n)$ time [15]. The following observation (which was also observed in [13]) leads to an $O(n^{3})$-time algorithm for the first task (solving SBSTR), which completes the analysis. We omit the basics on matroids, for which we refer the readers to [25].
Lemma 3.6 (cf. [13, Lemma 5]).
@@
Lemma 3.6 implies that SBST and SBSTR can be solved in polynomial time [^2] In contrast, it was recently shown that finding a strongly balanced spanning tree in a nonbipartite graph is NP-hard [4]. with the aid of matroid intersection algorithms. In particular, Brezovec, Cornuéjols, and Glover [5] proposed an $O(\tilde{n}\tilde{m}+\tilde{k}\tilde{n}^{2})$-time algorithm for the following special case: one matroid is a cycle matroid of a graph and the other is a partition matroid whose partition classes are induced by a stable set of the graph (i.e., each partition class is the set of edges incident to a vertex in the stable set), where $\tilde{n}$ and $\tilde{m}$ are the numbers of vertices and of edges in the graph, respectively, and $k$ is the size of the stable set inducing the partition classes. By the proof of Lemma 3.6, SBSTR completely fits this situation, and hence we can exactly solve it in $O(n^{3})$ time, where we have $\tilde{k}=|V^{+}|=n=\frac{\tilde{n}}{2}$ and $\tilde{m}\leq n^{2}$.
@@
Let $S_{\mathrm{opt}}$ be an optimal solution. We show $w(T\cup\overline{A_{\mathrm{out}}})=w(T)+w_{M}(A_{\mathrm{out}})\leq 2w(S_{\mathrm{opt}})$ (recall that $w_{M}\colon E(G_{M})\to\mathbb{R}_{\geq 0}$ is defined by (1)). The following two claims imply this.

### v3

@@ Abstract
Finding a minimum-weight strongly connected spanning subgraph of an edge-weighted directed graph is equivalent to the weighted version of the well-known strong connectivity augmentation problem. This problem is NP-hard, and a simple $2$-approximation algorithm was proposed by Frederickson and Jájá (1981); surprisingly, it still achieves the best known approximation ratio in general. Also, Bang-Jensen and Yeo (2008) showed that the unweighted problem is FPT (fixed-parameter tractable) parameterized by the difference from a trivial upper bound of the optimal value. In this paper, we consider a generalization related to the Dulmage–Mendelsohn decompositions of bipartite graphs instead of the strong connectivity of directed graphs, and extend these approximation and FPT results to the generalized setting.
@@ 1 Introduction
A directed graph is called strongly connected if there exists a path between any pair of vertices. The strong connectivity augmentation problem (SCA) is, given a directed graph and the weights of additional edges, to find a minimum-weight subset of additional edges making the graph strongly connected. While SCA is NP-hard in general, it can be solved in linear time in the unweighted case (when an edge between any pair of veritces can be added with the same weight) [10].
@@
The following problem, called the strongly connected spanning subgraph problem (SCSS), is polynomial-time equivalent to SCA in general: given a strongly connected directed graph with edge weight, find a minimum-weight spanning subgraph that is strongly connected. This problem is one of the most fundamental special cases of the survivable network design problem [27, 6]. Frederickson and Jájá [15] gave a $2$-approximation algorithm for SCSS; this is based on a very simple observation (see Section 3.1) and, surprisingly, it still achieves the best known approximation ratio in general. For the unweighted case (when all the edges have the same weight, which is already NP-hard in contrast to SCA), several algorithms with better approximation ratios [19, 28, 29, 18] have been proposed. Also, Bang-Jensen and Yeo [2] showed that the unweighted case is FPT parameterized by the difference from a trivial upper bound of the optimal value (see Sections 2.2 and 4.1); here, the problem is called FPT (fixed-parameter tractable) if it can be solved in $f(k)\cdot n^{O(1)}$ time, where $k$ and $f$ are the prescribed parameter and computable function, respectively, and $n$ is the input size.
@@
SCSS (or SCA) is generalized to several directions. One natural generalization is the $k$(-edge)-connected spanning subgraph (or $k$(-edge)-connectivity augmentation) problem, where a directed graph is called $k$-connected (or $k$-edge-connected) if there exists $k$ openly vertex-disjoint (or edge-disjoint) paths between any pair of vertices. The case when $k=1$ corresponds to SCSS (or SCA), and the general case is also well-studied [11, 13, 20, 25, 1].
@@
Bérczi, Iwata, Kato, and Yamaguchi [3] recently introduced another generalization of SCA, which is related to the Dulmage–Mendelsohn decomposition (the DM-decomposition) [8, 7] of a bipartite graph. The DM-decomposition gives a partition of the vertex set that reflects the structure of all maximum matchings therein, and a bipartite graph is DM-irreducible if its DM-decomposition consists of a single component. This property is closed under adding edges, and we analogously formulate the DM-irreducible spanning subgraph problem (DMISS): given a DM-irreducible bipartite graph with edge weight, find a minimum-weight spanning subgraph that is DM-irreducible. In terms of DMISS, Bérczi et al. [3] proposed polynomial-time algorithms for the case when the input bipartite graph is unbalanced (which reduces to the weighted matroid intersection problem), and for the case when the input graph is a balanced complete bipartite graph and the weight of each edge is 00 or $1$; here a bipartite graph is balanced if the two color classes have the same number of vertices. DMISS for balanced bipartite graphs generalizes SCSS (cf. [3, Section 2.3]), and the latter result extends the polynomial-time solvability of the unweighted case of SCA [10].
In this paper, we investigate DMISS for balanced bipartite graphs, whose algorithmic difficulty has remained widely open (other than the fact that it is NP-hard even in the unweighted case as so is SCSS). As natural but nontrivial results, by extending the approaches to SCSS by Frederickson and Jájá [15] and by Bang-Jensen and Yeo [2], we propose a $2$-approximation algorithm in general and an FPT algorithm for the unweighted case. On the approximation result, we first show that a naive extension of [15] gives a $3$-approximation algorithm, and then improve it to a $2$-approximation algorithm with the aid of the matroid intersection framework. On the FPT result, we demonstrate that the idea of [2] can be extended by carefully adjusting several key concepts and discussions.
@@
The rest of this paper is organized as follows. In Section 2, we describe necessary definitions, and formally state the problems and our results. In Section 3, we give a $2$-approximation algorithm for DMISS. In Section 4, we prove the unweighted case of DMISS is FPT parameterized by the difference from a trivial upper bound of the optimal value.
@@ 2.1 Definitions
For a graph $G$, we denote by $V(G)$ and $E(G)$ the vertex set and the edge set of $G$, respectively. A subgraph of $G$ is a graph $G^{\prime}$ such that $V(G^{\prime})\subseteq V(G)$, $E(G^{\prime})\subseteq E(G)$, and each edge $e\in E(G^{\prime})$ is between two vertices in $V(G^{\prime})$. A subgraph $G^{\prime}$ of $G$ is induced by a vertex set $X\subseteq V(G)$, denoted by $G[X]$, if $V(G^{\prime})=X$ and $E(G^{\prime})$ is the set of edges in $E(G)$ whose end vertices are both in $X$. We also define $G-X\coloneqq G[V(G)\setminus X]$ for $X\subseteq V(G)$. A subgraph $G^{\prime}$ of $G$ is spanning if $V(G^{\prime})=V(G)$. We often do not distinguish between a spanning subgraph $G^{\prime}$ and its edge set $E(G^{\prime})$. Also, we often specify a graph $G$ as the pair $(V,E)$ of its vertex set $V=V(G)$ and edge set $E=E(G)$.
Let $G=(V,E)$ be an undirected graph. A stable set in $G$ is a vertex set $X\subseteq V$ such that no edge in $E$ has its both end vertices in $X$. We say that $G$ is bipartite if there exists a bipartition $(V^{+},V^{-})$ of $V$ such that $V^{+}\cap V^{-}=\emptyset$, $V^{+}\cup V^{-}=V$, and both $V^{+}$ and $V^{-}$ are stable sets in $G$ (possibly $V^{+}$ or $V^{-}$ is empty). To specify the bipartition, we denote by $G=(V^{+},V^{-};E)$. For a vertex set $X$ on one side (i.e., $X\subseteq V^{+}$ or $X\subseteq V^{-}$), let $\Gamma_{G}(X)$ denote the neighborhood of $X$, i.e., the set of vertices adjacent to some vertex in $X$. A bipartite graph $G=(V^{+},V^{-};E)$ is balanced if $|V^{+}|=|V^{-}|$.
@@
The DM-decomposition [8, 7] of a bipartite graph is formally defined by the distributive lattice formed by the minimizers of a submodular function, and a bipartite graph is called DM-irreducible if its DM-decomposition consists of a single component. We here skip the formal definition and instead employ its characterization as an alternative “definition” (cf. [3, Section 2.2]). [^1] The equivalent condition is also called elementary, e.g., in [24, 12].
Lemma 2.1 (cf. [12, Theorem 2.4.18]).
@@
    $G$ has a perfect matching and $G_{M}$ is strongly connected for any perfect matching $M$ in $G$.
Suppose that a graph $G$ is associated with edge weight $w\colon E(G)\to\mathbb{R}$. We define the weight of each edge set $F\subseteq E(G)$ as $w(F)\coloneq\sum_{e\in F}w(e)$.
@@ 2.2 Problems and Our Results
We formally state the problems dealt with in this paper. Recall that DMISS for unbalanced bipartite graphs is polynomial-time solvable [3], which is not considered here.
@@
A strongly connected digraph $G$ with $w\colon E(G)\to\mathbb{R}_{\geq 0}$.
@@
A DM-irreducible balanced bipartite graph $G$ with $w\colon E(G)\to\mathbb{R}_{\geq 0}$.
@@
It is easy to observe that any inclusion-wise minimal strongly connected spanning subgraph of a digraph on $n$ vertices consists of at most $22n-$ 2 edges (by Lemma 3.1). Similarly, any inclusion-wise minimal DM-irreducible spanning subgraph of a balanced bipartite graph on $2n$ vertices ($n$ vertices on each side) consists of at most $23n-$ 2 edges (by Lemma 3.2). The following problems are the unweighted version of the above two problems parameterized by the difference from these trivial upper bounds; note that it is obvious from the definition that $|V(G)|$ edges are necessary in either problem.
@@
A strongly connected digraph $G$ on $n$ vertices.
@@
Test whether $G$ has a strongly connected spanning subgraph with at most $2n-2-k$ edges or not.
@@
A DM-irreducible balanced bipartite graph $G$ on $2n$ vertices.
@@
Test whether $G$ has a DM-irreducible spanning subgraph with at most $3n-2-k$ edges or not.
The results of this paper are summarized as follows, where $n=\frac{1}{2}|V(G)|$.
@@ 3 Approximation Algorithms for DMISS
In Section 3.1, as a base of our result, we review the $2$-approximation algorithm for SCSS proposed by Frederickson and Jájá [15]. In Section 3.2, we show a $3$-approximation algorithm for DMISS obtained by a naive extension of the result of Frederickson and Jájá. In Section 3.3, we improve it to a $2$-approximation algorithm with the aid of matroid intersection.
@@
We start with an elementary characterization of strongly connected digraphs.
@@
For the input $(G,w)$ of SCSS, the algorithm arbitrarily chooses a vertex $r$ and finds a minimum-weight spanning in-arborescence $A_{\mathrm{in}}$ and out-arborescence $A_{\mathrm{out}}$ rooted at $r$; the existence of $A_{\mathrm{in}}$ and $A_{\mathrm{out}}$ follows from Lemma 3.1 $(1\Rightarrow 3)$. This can be done in $O(m+n\log n)$ time [16], where $n=|V(G)|$ and $m=|E(G)|$. By Lemma 3.1 $(2\Rightarrow 1)$, the union of $A_{\mathrm{in}}$ and $A_{\mathrm{out}}$ results in a strongly connected spanning subgraph of $G$, and the algorithm outputs this. Since an optimal solution has a spanning in-arborescence and out-arborescence rooted at $r$ by Lemma 3.1 $(1\Rightarrow 3)$ and their weights are at least those of $A_{\mathrm{in}}$ and $A_{\mathrm{out}}$, respectively, the weight of the output solution is at most twice the optimal value.
@@
For the first task, polynomial-time exact algorithms initiated by the Hungarian method [23] are well-known. For the second task, we apply the $2$-approximation algorithm described in the previous section. We then output the union of the solutions (for the second solution, we take the set of corresponding edges in $G$). This intuitively leads to a $3$-approximation algorithm for DMISS, which can be formally shown as follows.
@@
We first give an easy but important observation.
@@
    $T$ has a perfect matching $M$ in $G$ and $\overrightarrow{T\setminus M}\cup\overleftarrow{M}$ is a spanning in-arborescence of $G_{M}$ (rooted at some vertex in $V^{+}$).
@@
Let us call a spanning tree satisfying Statement 2 in Lemma 3.5 strongly balanced, and say that the vertex $r$ is the root and the tree is rooted at $r$. By Lemma 3.2, every DM-irreducible balanced bipartite graph has a strongly balanced spanning tree, and then we formulate the first task as follows.
@@
A DM-irreducible balanced bipartite graph $G$ with $w\colon E(G)\to\mathbb{R}_{\geq 0}$.
@@
For the sake of simplicity of the following discussion, we consider the following problem, which is polynomial-time equivalent to the above problem and actually enough for our purpose.
Problem (SBST with Root (SBSTR)).
@@
A DM-irreducible balanced bipartite graph $G=(V^{+},V^{-};E)$ with $w\colon E\to\mathbb{R}_{\geq 0}$, and $r\in V^{+}$.
@@
Let $n=|V^{+}|=|V^{-}|$ and $m=|E|$. As we already mentioned, the second task (finding a minimum-weight spanning arborescence) can be done in $O(m+n\log n)$ time [16].
The following observation (which was also observed in [14]) leads to an $O(n^{3})$-time algorithm for the first task (solving SBSTR), which completes the analysis. We omit the basics on matroids, for which we refer the readers to [26].
Lemma 3.6 (cf. [14, Lemma 5]).
@@
Lemma 3.6 implies that SBST and SBSTR can be solved in polynomial time [^2] In contrast, it was recently shown that finding a strongly balanced spanning tree in a nonbipartite graph is NP-hard [4]. with the aid of matroid intersection algorithms. In particular, Brezovec, Cornuéjols, and Glover [5] proposed an $O(\tilde{n}\tilde{m}+\tilde{k}\tilde{n}^{2})$-time algorithm for the following special case: one matroid is the cycle matroid of a graph and the other is a partition matroid whose partition classes are induced by a stable set of the graph (i.e., each partition class is the set of edges incident to a vertex in the stable set), where $\tilde{n}$ and $\tilde{m}$ are the numbers of vertices and of edges in the graph, respectively, and $\tilde{k}$ is the size of the stable set inducing the partition classes. By the proof of Lemma 3.6, SBSTR completely fits this situation, and hence we can exactly solve it in $O(n^{3})$ time, where we have $\tilde{k}=n=\frac{\tilde{n}}{2}$ and $\tilde{m}\leq n^{2}$.
@@
Let $S_{\mathrm{opt}}$ be an optimal solution. We show $w(T\cup\overline{A_{\mathrm{out}}})=w(T)+w_{M}(A_{\mathrm{out}})\leq 2w(S_{\mathrm{opt}})$ (recall that $w_{M}\colon E(G_{M})\to\mathbb{R}_{\geq 0}$ is defined as (1)). The following two claims imply this.